OpenStudy (anonymous):

Let S be the part of the cone z^2 = x^2 + y^2 where z is between 1 and 2. (a) Write down the Monge parameterization of S and compute the surface element.

OpenStudy (anonymous):

not sure how to do the monge parameterization

OpenStudy (anonymous):

The parametrization is \[x(u, v) = <u, v, \pm \sqrt{u ^{2}+v ^{2} }>.\] For y > 0, we take the positive xquare root, and for y < 0, we take the negative square root.

OpenStudy (anonymous):

okay so u=x v=y and z=sqrt(u^2+v^2)

OpenStudy (anonymous):

I an uncertain about how to deal with the case y = 0. The parametrization as it stands now coners the lateral surface of the frustum except where the cone intersects y = 0., so it fails to completely cover that lateral surface. At least it is an open set like it's supposed to be.

OpenStudy (anonymous):

yeah well its close enough for me. thanks for the help!

OpenStudy (anonymous):

Cool beans:^)

OpenStudy (anonymous):

do you know how to Parameterize S using polar angle theta and z would it be 1<z<2 and 0<theta<2pi?

OpenStudy (anonymous):

In cylindrical coordinater, the cone has the equation, z = r. We disregard z = -r xince 1 < z <2. \[x = r \cos \theta ,y = r \sin \theta.\]

OpenStudy (anonymous):

The parametrization is now \[x(r,\theta) = <r \cos \theta, r \sin \theta, r)>.\]

OpenStudy (anonymous):

okay good thats what i ended up with. Im trying to find the bounds for the flux integral would theta be from 0 to 2pi and r and z are both from 1 to 2?

OpenStudy (anonymous):

They sure are.

OpenStudy (anonymous):

Great! thanks so much for the help :)

OpenStudy (anonymous):

:^)